# Classical and Quantum Heuristics for the Binary Paint Shop Problem

### v1

@@ V.1 Diminishing Returns of Level-1 RQAOA
Our large-scale benchmarks reveal a surprising and previously unreported trend: the performance of RQAOA1 deteriorates with problem size even when the QAOA1 subroutine is run with provably near-optimal parameters [82]. This stands in sharp contrast to prior guarantees showing that RQAOA1 can exactly solve unweighted cycles [30], complete graphs [75], and weighted bipartite graphs [76], as well as large-scale simulations on MaxCut and Ising models without external fields, which reported stable behaviour up to hundreds of qubits with no indication of degradation [30, 82]. Empirically, however, our kernel density estimates in fig. 3(a) shift monotonically toward higher paint-swap ratios as $n$ grows, providing the first clear evidence of RQAOA1 deteriorating at scale. Related work [82] also noted instability on Ising models with external fields, where RQAOA1 could perform extremely well on some instances but poorly on others.
@@ V.3 Beyond the Benchmarked Algorithms
In this work we restricted our benchmarks to classical heuristics with known guarantees for BPSP and shallow-depth QAOA variants previously shown to outperform the best classical algorithms on MaxCut. Other algorithms, while lacking proven performance for BPSP, could in principle achieve lower empirical paint-swap ratios. A natural candidate is the Goemans–Williamson (GW) algorithm [71]: since BPSP reduces to a weighted MaxCut instance, GW can be applied directly. However, two caveats arise. First, the presence of negative edge weights prevents the constant-factor approximation guarantee from carrying over. Second, GW requires solving a large semidefinite program whose memory demands grow rapidly with system size, often exceeding the capacity of standard hardware. For this reason we did not include GW in our large-scale comparisons; preliminary tests on small instances suggest that it can perform marginally better than XQAOA1, though its scalability is limited.
Beyond this, there exist many untested heuristics that could achieve lower empirical ratios on specific instances, including simple meta-heuristics such as randomised local search, albeit without provable guarantees. Finally, we emphasise that there is still considerable room between the conjectured asymptotic performance of XQAOA1, $\mathbb{E}_{\text{XQAOA}}[\Delta_{C}]\approx 0.357n$, and the recently established lower bound on the optimal solution, $\mathbb{E}_{\text{OPT}}[\Delta_{C}]\geq 0.214n-o(n)$  [50]. This gap suggests that further algorithmic improvements—classical, quantum, or hybrid—remain possible.
@@ B.2 Colouring Schemes
      $x^{(j)}_{k}=\begin{cases}x^{(j+1)}_{k},&k=1,2,\ldots,m_{j+1}-1,\\[2.0pt] x^{(j+1)}_{k+1},&k=m_{j+1},m_{j+1}+1,\ldots,2j.\end{cases}$ (35)
@@
        $g^{(j)}_{k}=\begin{cases}g^{(j-1)}_{k},&k=1,2,\ldots,m_{j}-1,\\[2.0pt] g^{(j-1)}_{k-1},&k=m_{j}+1,m_{j}+2,\ldots,2j-1.\end{cases}$ (36)
@@
    Reduction: Starting from $x^{(n)}$, peel off pairs from left to right; at each step note where the partner of the first car sits so the pair can be reinserted in the correct gap later. Formally, for $j=1,2,\ldots,n-1$:
@@
      $x^{(j)}_{k}=\begin{cases}x^{(j+1)}_{k+1},&k=1,2,\ldots,m_{j}-2,\\[2.0pt] x^{(j+1)}_{k+2},&k=m_{j}-1,m_{j},\ldots,2j,\end{cases}$ (37)
@@
        Embed the previously assigned symbols by copying the colouring from the shorter instance into $x^{(j)}$ at all positions except the two reinserted occurrences, which occupy indices $1$ and $m_{j}$:
        $h^{(j)}_{k}=\begin{cases}h^{(j-1)}_{k-1},&k=2,3,\ldots,m_{j}-1,\\[2.0pt] h^{(j-1)}_{k-2},&k=m_{j}+1,\ldots,2j.\end{cases}$ (38)

### v2

@@ V.1 Diminishing Returns of Level-1 RQAOA
Our large-scale benchmarks reveal a surprising and previously unreported trend: the performance of RQAOA1 deteriorates with problem size even when the QAOA1 subroutine is run with provably near-optimal parameters [vijendran2025near]. This stands in sharp contrast to prior guarantees showing that RQAOA1 can exactly solve unweighted cycles [bravyi2020obstacles], complete graphs [bae2024recursive], and weighted bipartite graphs [bae2024modified], as well as large-scale simulations on MaxCut and Ising models without external fields, which reported stable behaviour up to hundreds of qubits with no indication of degradation [bravyi2020obstacles, vijendran2025near]. Empirically, however, our kernel density estimates in fig. 3(a) shift monotonically toward higher paint-swap ratios as $n$ grows, providing the first clear evidence of RQAOA1 deteriorating at scale. Related work [vijendran2025near] also noted instability on Ising models with external fields, where RQAOA1 could perform extremely well on some instances but poorly on others.
@@ V.3 Beyond the Benchmarked Algorithms
In this work we restricted our benchmarks to classical heuristics with known guarantees for BPSP and shallow-depth QAOA variants previously shown to outperform the best classical algorithms on MaxCut. Other algorithms, while lacking proven performance for BPSP, could in principle achieve lower empirical paint-swap ratios. A natural candidate is the Goemans–Williamson (GW) algorithm [goemans1995improved]: since BPSP reduces to a weighted MaxCut instance, GW can be applied directly. However, two caveats arise. First, the presence of negative edge weights prevents the constant-factor approximation guarantee from carrying over. Second, GW requires solving a large semidefinite program whose memory demands grow rapidly with system size, often exceeding the capacity of standard hardware. For this reason we did not include GW in our large-scale comparisons; preliminary tests on small instances suggest that it can perform marginally better than XQAOA1, though its scalability is limited.
Beyond this, there exist many untested heuristics that could achieve lower empirical ratios on specific instances, including simple meta-heuristics such as randomised local search, albeit without provable guarantees. Finally, we emphasise that there is still considerable room between the conjectured asymptotic performance of XQAOA1, $\mathbb{E}_{\text{XQAOA}}[\Delta_{C}]\approx 0.357n$, and the recently established lower bound on the optimal solution, $\mathbb{E}_{\text{OPT}}[\Delta_{C}]\geq 0.214n-o(n)$  [hanvcl2023improved]. This gap suggests that further algorithmic improvements—classical, quantum, or hybrid—remain possible.
@@ B.2 Colouring Schemes
      $x^{(j)}_{k}=\begin{cases}x^{(j+1)}_{k},&k=1,2,\ldots,m_{j+1}-1,\\[2.0pt] x^{(j+1)}_{k+1},&k=m_{j+1},m_{j+1}+1,\ldots,2j.\end{cases}$
@@
        $g^{(j)}_{k}=\begin{cases}g^{(j-1)}_{k},&k=1,2,\ldots,m_{j}-1,\\[2.0pt] g^{(j-1)}_{k-1},&k=m_{j}+1,m_{j}+2,\ldots,2j-1.\end{cases}$
@@
    Reduction: Starting from $x^{(n)}$, peel off pairs from left to right; at each step note where the partner of the first car sits so the pair can be reinserted in the correct gap later. Formally, for $j=n-1,n-2,\ldots,1$:
@@
      $x^{(j)}_{k}=\begin{cases}x^{(j+1)}_{k+1},&k=1,2,\ldots,m_{j}-2,\\[2.0pt] x^{(j+1)}_{k+2},&k=m_{j}-1,m_{j},\ldots,2j,\end{cases}$
@@
        Embed the previously assigned symbols by copying the colouring from the shorter instance into $x^{(j)}$ at all positions except the two reinserted occurrences, which occupy indices $1$ and $m_{j-1}$:
        $h^{(j)}_{k}=\begin{cases}h^{(j-1)}_{k-1},&k=2,3,\ldots,m_{j-1}-1,\\[2.0pt] h^{(j-1)}_{k-2},&k=m_{j-1}+1,\ldots,2j.\end{cases}$